### v1

# Structured Pruning for Diverse Best-of- Reasoning Optimization

Abstract
Model pruning in transformer-based language models, traditionally viewed as a means of achieving computational savings, can enhance the modelās reasoning capabilities. In this work, we uncover a surprising phenomenon: the selective pruning of certain attention heads leads to improvements in reasoning performance, particularly on challenging tasks. Motivated by this observation, we propose SPRINT, a novel contrastive learning framework that dynamically selects the optimal head and layer to prune during inference. By aligning question embeddings with head embeddings, SPRINT identifies those pruned-head configurations that result in more accurate reasoning. Extensive experiments demonstrate that our method significantly outperforms traditional best-of- and random head selection strategies on the MATH500 and GSM8K datasets.
Structured Pruning for Diverse Best-of- Reasoning Optimization
Hieu Trung Nguyen, āāBao Nguyen, āāViet Anh Nguyen The Chinese University of Hong Kong {thnguyen, nbnguyen, nguyen}@se.cuhk.edu.hk
1 Introduction
Language models (LMs) have recently made remarkable progress in understanding and reasoning with human language, attracting considerable research attentionĀ Brown etĀ al. (2020); Hoffmann etĀ al. (2022); Chowdhery etĀ al. (2022); Rae etĀ al. (2021); Raffel etĀ al. (2019). This unprecedented evolution opened up a wide range of practical applications, including content generationĀ Li etĀ al. (2024a); Liu etĀ al. (2023b); Achiam etĀ al. (2023), virtual assistanceĀ Sezgin (2024); Doe etĀ al. (2023); Garcia etĀ al. (2023), population simulationĀ Bui etĀ al. (2025), and problem solvingĀ Ahn etĀ al. (2024); Imani etĀ al. (2023); Lewkowycz etĀ al. (2022).
Despite these advances, the reliability and correctness of LM-generated answers remain a significant concern. Analyses of their output frequently reveal faulty reasoning and factual inaccuraciesĀ Ji etĀ al. (2023); Xu etĀ al. (2024); Nguyen etĀ al. (2025a); Jiang etĀ al. (2025); Li etĀ al. (2024b). This issue is particularly pronounced in tasks that require advanced reasoning, such as automated theorem provingĀ Wu etĀ al. (2022), mathematical problem solvingĀ Trinh etĀ al. (2024), or heuristic discoveryĀ Romera-Paredes etĀ al. (2024). LLMs often struggle to produce accurate responses in these scenarios in a single pass.
To address this challenge, iterative generation strategies have been employed to refine and select the most appropriate response, often by combining language generation with aggregation or search techniques such as best-of-N samplingĀ Stiennon etĀ al. (2020) or Monte Carlo Tree SearchĀ Xie etĀ al. (2024). Throughout this paper, we refer to these iterative approaches as reasoning methods.
Current reasoning methods typically leverage the diversity induced by the LLMās decoding process through parameters such as temperature, top-, and top- to generate multiple candidate answersĀ Stiennon etĀ al. (2020); Xie etĀ al. (2024). Although this technique introduces textual diversity and paraphrasing, it does not capture fundamentally different viewpoints, since all candidates are produced by the same underlying model with similar reasoning capabilities. One potential remedy is to employ a variety of LMs to generate answers; however, this approach is often computationally expensive and memory inefficient.
At the same time, recent research on model pruning has shown that comparable performance can be achieved by selectively disabling certain heads in the transformer architectureĀ Ma etĀ al. (2023); Wang etĀ al. (2019). Removing an attention head does not necessarily weaken the overall ability of the model because a head may introduce noise into the decision-making process, leading to incorrect answers. In fact, our empirical observations suggest that pruning certain heads can improve the reasoning methodsā performance on specific categories of questions; SectionĀ 3.2 discusses this phenomenon in more detail. This finding motivates our approach: we aim to construct a pool of pruned models from one base model with diverse reasoning capabilities without incurring additional computational or memory overhead.
Contributions. We begin by analyzing and providing new insights into the surprising phenomenon that pruning attention heads can improve the performance of the language model on the MATH benchmark. Motivated by these insights, we introduce the Structured PRunIng for diverSe reasoNing opTimization (SPRINT) framework. SPRINT first learns a set of embeddings for each layer-attention head pair using contrastive learning, guided by a novel Diversity-Promoted Contrastive Loss. This loss function encourages the question embedding to move closer to the embeddings of heads whose pruning leads to correct answers for that question. Finally, we leverage SPRINT to enhance best-of- sampling by pruning the top nearest layer-model configurations, determined by the distance between the learned embeddings and the question embedding.
Experiments on the MATH500 and GSM8K datasets show that our method outperforms traditional best-of- sampling approaches, achieving higher accuracy without introducing inference-time overhead.
Our paper unfolds as follows: SectionĀ 2 discusses related work on reasoning in LMs. SectionĀ 3 investigates the impact of pruning individual attention heads in transformer models, revealing that certain heads can be removed without performance loss and, in some cases, even improve accuracy. SectionĀ 4 delineates our SPRINT framework for matching heads for pruning with input questions, and SectionĀ 5 presents the extensive numerical results of the mathematical reasoning task.
2 Related Works
A straightforward decoding process may struggle to produce accurate solutions for complex reasoning tasks. To address this, Self-Consistency (SC) improves reliability by generating multiple outputs and selecting the final response by majority votingĀ Wang etĀ al. (2022). A similar approach, best-of-N sampling, employs a reward model or function to select the response with the highest reward Stiennon etĀ al. (2020). While both techniques enhance output quality, they also increase computational cost proportional to the number of samples.
To explore reasoning pathways more effectively, tree-search-based methods have been introduced, such as Tree-of-Thought Yao etĀ al. (2024), Monte Carlo Tree Search (MCTS) Feng etĀ al. (2023); Zhang etĀ al. (2024); Guan etĀ al. (2025), and Forest-of-Thought Bi etĀ al. (2024). Damani etĀ al. (2024) highlights that tree-based search strategies are more effective in discovering correct solutions than parallel sampling, particularly for complex problems.
Beyond reasoning strategies, scaling inference-time computation also plays a crucial role in improving output quality. Beeching etĀ al. (2024) shows that the accuracy on the MATH-500 benchmark improves as the test-time computation (that is, the number of generations per problem) increases for methods such as best-of-N, beam search, and diverse verifier tree search (DVTS). Similarly, Guan etĀ al. (2025) achieved an average accuracy of 53.3% on the USA Math Olympiad (AIME) by conducting extensive MCTS rollouts.
Our work is related to recent research on interventions in the internal state of language modelsĀ Wu etĀ al. (2024); Li etĀ al. (2023); Nguyen etĀ al. (2025b); Liu etĀ al. (2023a); Nguyen etĀ al. (2025c). Although prior approaches often focus on learning intervention vectors to add to model activations, instead we investigate the effect of excluding specific components of the internal state of the model.
3 Does Pruning an Attention Head Improve LMās Performance?
We consider a transformer-based model, where each layer contains attention heads. In each layer , the outputs of the attention heads are denoted as , where each . The concatenated vector, with , is then passed to a linear layer to produce the output of a multi-head attention block.
Attention head pruning. In our study, we define head pruning as zeroing out the output vector of a specific attention head before the concatenation step, as illustrated in FigureĀ 1. In modern decoder-only architecturesĀ , this is equivalent to pruning the input of layer.
In this section, we detail our experimental procedure for attention head pruning and present a surprising empirical phenomenon that motivates our subsequent work.
3.1 Experimental Procedure
To systematically examine the effect of individual head pruning, we follow the procedure below:
1.
Layer and Head Selection: We select layers from the model, yielding a total of attention heads. For each head, we create a variant of the base model with that specific head pruned, resulting in model variants in addition to the base model. We denote this model pool as .
2.
Measuring the Impact: For each model , we generate solutions on the MATH500 dataset, which consists of samples. The predictions are represented by a binary accuracy vector , where if sample is solved correctly and otherwise. Evaluating all models yields a matrix of accuracy values.
In this experiment, we study two models specialized for mathematical reasoning, Qwen2.5-Math-1.5B-Instruct and Qwen2.5-Math-7B-Instruct, as well as two general-purpose models, Qwen2.5-7B-Instruct, and Meta-Llama-3-8B-Instruct. Our experiments involve evaluating model performance with and without head pruning on 500 samples from the MATH500 datasetĀ Hendrycks etĀ al. (2021).
3.2 Empirical Observations
Our empirical results, shown in TablesĀ 2, 3, 4, and 1, reveal several important and sometimes surprising findings about attention head pruning. Contrary to the common belief that all attention heads are essential for optimal model performance, we find that pruning individual heads often does not reduce accuracy, and in many cases, actually increases it. The impact of pruning depends strongly on both the specific head and its layer: while some heads are crucial (with pruning causing sharp drops in accuracy), many others are unnecessary, and removing them can leave performance unchanged or even lead to notable improvements. For instance, in the āCounting & Probabilityā category, pruning head 0 in layer 15 of Qwen2.5-Math-1.5B-Instruct increased accuracy by 16% over the unpruned baseline.
We further illustrate this trend in FigureĀ 2, which displays violin plots of additive performance gains (the difference between the best pruned head accuracy and the baseline accuracy) across all subcategories for each model. These plots show that head pruning consistently produces positive gains for all evaluated models. The specialized mathematical models (Qwen2.5-Math-1.5B-Instruct and Qwen2.5-Math-7B-Instruct) exhibit higher median gains and more frequent large improvements, particularly in domains such as āCounting & Probabilityā and āGeometryā. In contrast, the general-purpose models (Qwen2.5-7B-Instruct and Meta-Llama-3-8B-Instruct) show more modest but still positive and consistent benefits from pruning.
Overall, these patterns suggest that some attention heads contribute redundant or even harmful information, and that pruning can act as a useful form of regularization. Our findings highlight the significant redundancy present in multi-head attention and indicate promising directions for simplifying models and improving efficiency through targeted head pruning.
4 Methodology
Motivated by these counterintuitive findings, we propose a framework called Structured Pruning for Diverse Reasoning Optimization (SPRINT) to leverage selective head pruning to improve reasoning performance. At the core of SPRINT is a contrastive learning approach that aligns question embeddings with attention head embeddings to identify the heads most beneficial to prune. Initially, every head is associated with a learnable vector from a set , where each is randomly initialized. A sentence embedding model, denoted by , processes each question , after which a linear transformation produces the question embedding . For each question, we define the set of pruned-head models as and partition this set into a positive subset containing heads that result in correct answers to question , and a negative subset .
To further quantify the correspondence between different pruned-head models, we introduce a similarity measure between any two heads and as follows:
which computes the fraction of questions for which both heads produce the same outcome. We jointly optimize the transformation and the head embeddings with the following loss function:
Our joint optimization loss consists of two components. The first term aims to align the question embedding with the embeddings of those pruned-head models that yield the correct answer. The loss employs a softmax formulation for each question, over the negative squared Euclidean distances between and the head embeddings. This loss encourages the model to produce a lower distance, and thus a higher similarity, for those heads in the positive set , which results in a correct answer.
The second term of the loss addresses the need for diversity among head embeddings. By summing the pairwise distances weighted by the similarity measure , it acts as a regularizer that encourages distinctiveness between head embeddings, especially for those heads that tend to yield similar outcomes. This term effectively pushes similar heads apart in the embedding space to reduce redundancy and promote a richer, more discriminative representation. Overall, combining these two terms helps align the question representation with high-performing pruned models while maintaining a diverse set of head embeddings.
Inference. At inference time, we obtain the embedding for the new question by feedforwarding . The optimal head to prune for a given test question is identified by selecting the head whose embedding minimizes the squared Euclidean distance to the question embedding, i.e.,
We integrate best-of- reasoning into our framework to enhance the performance by selecting the top- pruned-head configurations based on their proximity to . We obtain candidate answers via greedy decoding with a temperature of 0.
Overall, our unified methodology demonstrates that pruning the attention heads, guided by a contrastive alignment strategy, can improve model performance. Thus, the SPRINT framework identifies which heads benefit from pruning and supports a dynamic simulation approach to generate optimal predictions.
5 Numerical Experiments
Datasets. For our experiments, we use the MATH and GSM8K datasets. We provide more details in the appendix.
Base Models. As a proof of concept, we select Qwen2.5-Math-1.5B-Instruct as our generation model. This choice is motivated by its strong empirical gains observed in FigureĀ 2. All experiments are performed using the original, unquantized version of this model. We consider five layers {5, 10, 15, 20, 25} for pruning.
Performance metrics. To assess performance, we use the Pass@ metric. Specifically, each method is used to sample candidate answers, and the accuracy is computed based on the best answer as determined by an oracle.
Baselines. We compare our methods against two baselines: multinomial sampling, commonly used for Best-of- Reasoning, and a random head selection strategy. We perform Best-of- evaluations using three temperature values for multinomial sampling: 0.1, 0.6, and 0.9. For the random head selection strategy, we use a greedy approach to identify the heads that solve the most samples in the training set. During testing, we randomly select heads and measure Pass@.
The projectās repository is available at https://github.com/HieuNT91/attention_pruning.
5.1 Quantitative Results
FigureĀ 4 summarizes our primary findings. Our method consistently outperforms the baselines for nearly all values of across the MATH and GSM8K datasets. Notably, the greatest improvements are observed at lower values , while the performance gains become more gradual as increases.
6 Conclusion
In this work, we demonstrated that pruning attention heads, a technique traditionally viewed as harmful, can enhance the reasoning capabilities of transformer-based language models. Our study revealed that the selective removal of specific heads mitigates the propagation of redundant or noisy signals and enhances overall model performance. We developed a contrastive learning framework to harness this phenomenon that dynamically selects the optimal head and layer to prune at inference time. Experimental results confirm that our approach significantly outperforms traditional Best-of- methods and random head selection strategies.
Limitations
Due to the space limitations of a short paper, our work does not experiment with a more diverse range of model architectures and sizes. Our work primarily focuses on the mathematical domain, but it can be extended to other scientific domains that require LMsā reasoning.
Appendix A Implementation details.
Datasets. The MATH dataset contains 12,500 math problems, with 12,000 problems designated for training and 500 for testing. Similarly, the GSM8K dataset comprises 8,792 math problems, with 7,473 used for training and 1,319 for testing. In our experiments, we subsample 1,500 training examples from the respective training sets. The head embeddings are then trained using the subsampled training set. Finally, we evaluate performance on the test splits of the aforementioned datasets.
Appendix B Empirical Observation on the Impact of Attention Head Pruning.
FigureĀ 5 shows that pruning an attention head in Qwen-2.5-1.5B Math Instruct model mostly leads to improvement in accuracy@1 on the MATH500 dataset.
| Algebra | Counting & Probability | Geometry | Intermediate Algebra | Number Theory | Prealgebra | Precalculus | ||||||||||||||||||||||
| Layer | 5 | 10 | 15 | 31 | 5 | 10 | 15 | 31 | 5 | 10 | 15 | 31 | 5 | 10 | 15 | 31 | 5 | 10 | 15 | 31 | 5 | 10 | 15 | 31 | 5 | 10 | 15 | 31 |
| Prune Head 0 | 0.35 | 0.37 | 0.32 | 0.32 | 0.11 | 0.11 | 0.13 | 0.16 | 0.2 | 0.24 | 0.2 | 0.22 | 0.13 | 0.14 | 0.14 | 0.16 | 0.19 | 0.19 | 0.21 | 0.18 | 0.4 | 0.38 | 0.41 | 0.39 | 0.14 | 0.12 | 0.14 | 0.12 |
| Prune Head 1 | 0.31 | 0.37 | 0.32 | 0.34 | 0.11 | 0.11 | 0.08 | 0.08 | 0.2 | 0.22 | 0.22 | 0.24 | 0.14 | 0.13 | 0.14 | 0.13 | 0.19 | 0.24 | 0.19 | 0.24 | 0.41 | 0.37 | 0.37 | 0.4 | 0.11 | 0.14 | 0.11 | 0.11 |
| Prune Head 2 | 0.33 | 0.32 | 0.32 | 0.34 | 0.16 | 0.13 | 0.11 | 0.11 | 0.22 | 0.15 | 0.22 | 0.2 | 0.14 | 0.13 | 0.15 | 0.15 | 0.27 | 0.18 | 0.24 | 0.18 | 0.34 | 0.34 | 0.41 | 0.41 | 0.11 | 0.09 | 0.09 | 0.12 |
| Prune Head 3 | 0.35 | 0.31 | 0.34 | 0.33 | 0.08 | 0.11 | 0.11 | 0.13 | 0.17 | 0.22 | 0.24 | 0.22 | 0.13 | 0.1 | 0.16 | 0.16 | 0.19 | 0.16 | 0.21 | 0.19 | 0.38 | 0.48 | 0.35 | 0.34 | 0.09 | 0.18 | 0.11 | 0.09 |
| Prune Head 4 | 0.31 | 0.34 | 0.3 | 0.33 | 0.11 | 0.11 | 0.16 | 0.11 | 0.22 | 0.22 | 0.22 | 0.22 | 0.14 | 0.13 | 0.15 | 0.14 | 0.15 | 0.21 | 0.23 | 0.19 | 0.45 | 0.39 | 0.44 | 0.38 | 0.12 | 0.07 | 0.11 | 0.11 |
| Prune Head 5 | 0.33 | 0.33 | 0.36 | 0.33 | 0.16 | 0.08 | 0.13 | 0.11 | 0.2 | 0.24 | 0.24 | 0.24 | 0.12 | 0.12 | 0.1 | 0.11 | 0.16 | 0.27 | 0.23 | 0.21 | 0.39 | 0.37 | 0.4 | 0.43 | 0.11 | 0.12 | 0.11 | 0.09 |
| Prune Head 6 | 0.35 | 0.35 | 0.38 | 0.32 | 0.13 | 0.11 | 0.13 | 0.13 | 0.2 | 0.2 | 0.22 | 0.24 | 0.1 | 0.13 | 0.1 | 0.14 | 0.26 | 0.18 | 0.23 | 0.19 | 0.38 | 0.43 | 0.39 | 0.39 | 0.09 | 0.18 | 0.12 | 0.11 |
| Prune Head 7 | 0.35 | 0.34 | 0.31 | 0.31 | 0.11 | 0.05 | 0.13 | 0.13 | 0.17 | 0.24 | 0.17 | 0.22 | 0.11 | 0.11 | 0.12 | 0.16 | 0.21 | 0.21 | 0.21 | 0.19 | 0.39 | 0.4 | 0.43 | 0.35 | 0.07 | 0.09 | 0.14 | 0.11 |
| Prune Head 8 | 0.35 | 0.37 | 0.31 | 0.34 | 0.13 | 0.18 | 0.11 | 0.16 | 0.22 | 0.24 | 0.22 | 0.2 | 0.11 | 0.12 | 0.12 | 0.13 | 0.23 | 0.21 | 0.13 | 0.21 | 0.41 | 0.38 | 0.34 | 0.4 | 0.12 | 0.12 | 0.18 | 0.11 |
| Prune Head 9 | 0.35 | 0.31 | 0.36 | 0.32 | 0.13 | 0.08 | 0.08 | 0.13 | 0.22 | 0.17 | 0.24 | 0.22 | 0.13 | 0.1 | 0.13 | 0.16 | 0.21 | 0.23 | 0.26 | 0.21 | 0.41 | 0.39 | 0.34 | 0.39 | 0.07 | 0.11 | 0.16 | 0.11 |
| Prune Head 10 | 0.33 | 0.37 | 0.38 | 0.34 | 0.16 | 0.08 | 0.08 | 0.13 | 0.24 | 0.2 | 0.17 | 0.24 | 0.14 | 0.09 | 0.14 | 0.15 | 0.19 | 0.21 | 0.21 | 0.21 | 0.4 | 0.4 | 0.39 | 0.4 | 0.11 | 0.18 | 0.12 | 0.12 |
| Prune Head 11 | 0.33 | 0.34 | 0.32 | 0.33 | 0.11 | 0.05 | 0.18 | 0.13 | 0.22 | 0.24 | 0.27 | 0.2 | 0.16 | 0.12 | 0.13 | 0.16 | 0.21 | 0.18 | 0.18 | 0.19 | 0.37 | 0.38 | 0.37 | 0.37 | 0.14 | 0.12 | 0.12 | 0.11 |
| Prune Head 12 | 0.35 | 0.34 | 0.33 | 0.35 | 0.11 | 0.13 | 0.18 | 0.13 | 0.27 | 0.22 | 0.2 | 0.24 | 0.14 | 0.12 | 0.14 | 0.12 | 0.24 | 0.19 | 0.19 | 0.21 | 0.38 | 0.44 | 0.43 | 0.39 | 0.12 | 0.12 | 0.12 | 0.07 |
| Prune Head 13 | 0.35 | 0.35 | 0.27 | 0.35 | 0.13 | 0.05 | 0.18 | 0.18 | 0.15 | 0.24 | 0.1 | 0.2 | 0.16 | 0.12 | 0.11 | 0.14 | 0.23 | 0.21 | 0.16 | 0.21 | 0.41 | 0.41 | 0.28 | 0.43 | 0.11 | 0.12 | 0.12 | 0.11 |
| Prune Head 14 | 0.35 | 0.38 | 0.32 | 0.27 | 0.11 | 0.24 | 0.11 | 0.08 | 0.2 | 0.2 | 0.17 | 0.17 | 0.14 | 0.13 | 0.13 | 0.12 | 0.23 | 0.23 | 0.16 | 0.18 | 0.38 | 0.38 | 0.3 | 0.37 | 0.12 | 0.12 | 0.11 | 0.14 |
| Prune Head 15 | 0.3 | 0.31 | 0.35 | 0.32 | 0.13 | 0.11 | 0.16 | 0.16 | 0.15 | 0.17 | 0.2 | 0.22 | 0.1 | 0.16 | 0.09 | 0.13 | 0.23 | 0.21 | 0.23 | 0.19 | 0.41 | 0.35 | 0.35 | 0.41 | 0.14 | 0.11 | 0.11 | 0.07 |
| Prune Head 16 | 0.35 | 0.35 | 0.33 | 0.33 | 0.13 | 0.11 | 0.13 | 0.13 | 0.22 | 0.24 | 0.22 | 0.24 | 0.13 | 0.14 | 0.15 | 0.15 | 0.18 | 0.18 | 0.21 | 0.21 | 0.39 | 0.4 | 0.44 | 0.39 | 0.18 | 0.11 | 0.12 | 0.11 |
| Prune Head 17 | 0.33 | 0.36 | 0.35 | 0.35 | 0.13 | 0.13 | 0.13 | 0.11 | 0.17 | 0.2 | 0.24 | 0.22 | 0.12 | 0.11 | 0.14 | 0.15 | 0.13 | 0.18 | 0.21 | 0.16 | 0.34 | 0.37 | 0.39 | 0.4 | 0.12 | 0.12 | 0.12 | 0.11 |
| Prune Head 18 | 0.35 | 0.34 | 0.36 | 0.36 | 0.11 | 0.13 | 0.05 | 0.08 | 0.2 | 0.2 | 0.29 | 0.17 | 0.14 | 0.12 | 0.11 | 0.15 | 0.24 | 0.24 | 0.23 | 0.23 | 0.35 | 0.37 | 0.41 | 0.37 | 0.16 | 0.14 | 0.11 | 0.11 |
| Prune Head 19 | 0.32 | 0.36 | 0.34 | 0.33 | 0.11 | 0.13 | 0.13 | 0.13 | 0.22 | 0.22 | 0.2 | 0.27 | 0.1 | 0.12 | 0.11 | 0.15 | 0.16 | 0.21 | 0.19 | 0.19 | 0.35 | 0.37 | 0.39 | 0.35 | 0.07 | 0.16 | 0.12 | 0.11 |
| Prune Head 20 | 0.31 | 0.37 | 0.34 | 0.35 | 0.13 | 0.11 | 0.05 | 0.16 | 0.22 | 0.24 | 0.22 | 0.2 | 0.12 | 0.14 | 0.12 | 0.15 | 0.21 | 0.21 | 0.16 | 0.21 | 0.38 | 0.37 | 0.37 | 0.38 | 0.11 | 0.09 | 0.11 | 0.12 |
| Prune Head 21 | 0.32 | 0.32 | 0.33 | 0.34 | 0.13 | 0.13 | 0.16 | 0.11 | 0.24 | 0.15 | 0.24 | 0.24 | 0.14 | 0.09 | 0.16 | 0.15 | 0.23 | 0.11 | 0.16 | 0.21 | 0.4 | 0.38 | 0.39 | 0.38 | 0.11 | 0.14 | 0.11 | 0.11 |
| Prune Head 22 | 0.35 | 0.37 | 0.35 | 0.35 | 0.05 | 0.13 | 0.16 | 0.08 | 0.22 | 0.2 | 0.2 | 0.24 | 0.14 | 0.14 | 0.16 | 0.15 | 0.15 | 0.19 | 0.21 | 0.21 | 0.38 | 0.4 | 0.39 | 0.39 | 0.12 | 0.14 | 0.14 | 0.11 |
| Prune Head 23 | 0.35 | 0.40 | 0.33 | 0.35 | 0.13 | 0.13 | 0.11 | 0.16 | 0.2 | 0.22 | 0.22 | 0.22 | 0.15 | 0.13 | 0.13 | 0.18 | 0.23 | 0.18 | 0.23 | 0.23 | 0.39 | 0.38 | 0.41 | 0.4 | 0.11 | 0.14 | 0.12 | 0.11 |
| Prune Head 24 | 0.36 | 0.33 | 0.39 | 0.32 | 0.21 | 0.13 | 0.05 | 0.08 | 0.29 | 0.27 | 0.2 | 0.1 | 0.12 | 0.14 | 0.14 | 0.13 | 0.18 | 0.24 | 0.18 | 0.29 | 0.38 | 0.35 | 0.4 | 0.41 | 0.11 | 0.11 | 0.09 | 0.12 |
| Prune Head 25 | 0.34 | 0.32 | 0.33 | 0.34 | 0.11 | 0.13 | 0.08 | 0.11 | 0.22 | 0.2 | 0.2 | 0.2 | 0.09 | 0.13 | 0.12 | 0.14 | 0.21 | 0.21 | 0.18 | 0.21 | 0.38 | 0.44 | 0.46 | 0.43 | 0.11 | 0.11 | 0.12 | 0.11 |
| Prune Head 26 | 0.32 | 0.27 | 0.35 | 0.40 | 0.11 | 0.16 | 0.13 | 0.08 | 0.2 | 0.24 | 0.2 | 0.17 | 0.14 | 0.12 | 0.12 | 0.14 | 0.18 | 0.18 | 0.16 | 0.23 | 0.4 | 0.41 | 0.39 | 0.4 | 0.12 | 0.14 | 0.14 | 0.12 |
| Prune Head 27 | 0.34 | 0.31 | 0.33 | 0.34 | 0.08 | 0.11 | 0.08 | 0.13 | 0.22 | 0.29 | 0.24 | 0.27 | 0.14 | 0.14 | 0.16 | 0.14 | 0.19 | 0.26 | 0.18 | 0.26 | 0.38 | 0.35 | 0.41 | 0.39 | 0.11 | 0.09 | 0.09 | 0.07 |
| Prune Head 28 | 0.32 | 0.31 | 0.35 | 0.35 | 0.11 | 0.21 | 0.13 | 0.13 | 0.2 | 0.22 | 0.24 | 0.2 | 0.11 | 0.1 | 0.15 | 0.14 | 0.21 | 0.19 | 0.16 | 0.21 | 0.38 | 0.4 | 0.39 | 0.38 | 0.12 | 0.12 | 0.11 | 0.09 |
| Prune Head 29 | 0.37 | 0.38 | 0.34 | 0.33 | 0.13 | 0.13 | 0.08 | 0.11 | 0.29 | 0.2 | 0.22 | 0.24 | 0.14 | 0.15 | 0.15 | 0.15 | 0.21 | 0.16 | 0.19 | 0.21 | 0.41 | 0.43 | 0.39 | 0.4 | 0.07 | 0.14 | 0.14 | 0.11 |
| Prune Head 30 | 0.35 | 0.34 | 0.35 | 0.33 | 0.13 | 0.16 | 0.16 | 0.05 | 0.2 | 0.17 | 0.24 | 0.24 | 0.13 | 0.13 | 0.14 | 0.15 | 0.19 | 0.23 | 0.19 | 0.21 | 0.39 | 0.38 | 0.37 | 0.38 | 0.11 | 0.14 | 0.16 | 0.11 |
| Prune Head 31 | 0.31 | 0.33 | 0.37 | 0.33 | 0.13 | 0.13 | 0.16 | 0.16 | 0.27 | 0.24 | 0.2 | 0.24 | 0.14 | 0.12 | 0.13 | 0.14 | 0.19 | 0.16 | 0.23 | 0.19 | 0.4 | 0.4 | 0.39 | 0.39 | 0.11 | 0.14 | 0.11 | 0.11 |
| No Prune | 0.33 | 0.33 | 0.33 | 0.33 | 0.11 | 0.11 | 0.11 | 0.11 | 0.24 | 0.24 | 0.24 | 0.24 | 0.15 | 0.15 | 0.15 | 0.15 | 0.21 | 0.21 | 0.21 | 0.21 | 0.39 | 0.39 | 0.39 | 0.39 | 0.11 | 0.11 | 0.11 | 0.11 |
| Algebra | Counting & Probability | Geometry | Intermediate Algebra | Number Theory | Prealgebra | Precalculus | ||||||||||||||||||||||
| Layer | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 |
| Prune Head 0 | 0.92 | 0.9 | 0.91 | 0.91 | 0.63 | 0.63 | 0.79 | 0.66 | 0.56 | 0.49 | 0.54 | 0.51 | 0.59 | 0.62 | 0.55 | 0.62 | 0.82 | 0.87 | 0.89 | 0.87 | 0.82 | 0.82 | 0.79 | 0.78 | 0.62 | 0.59 | 0.57 | 0.68 |
| Prune Head 1 | 0.92 | 0.92 | 0.92 | 0.92 | 0.71 | 0.68 | 0.63 | 0.63 | 0.51 | 0.56 | 0.59 | 0.51 | 0.6 | 0.52 | 0.53 | 0.58 | 0.84 | 0.85 | 0.85 | 0.90 | 0.82 | 0.77 | 0.78 | 0.79 | 0.62 | 0.62 | 0.64 | 0.66 |
| Prune Head 2 | 0.93 | 0.91 | 0.91 | 0.91 | 0.74 | 0.66 | 0.66 | 0.71 | 0.59 | 0.54 | 0.63 | 0.54 | 0.61 | 0.58 | 0.54 | 0.59 | 0.92 | 0.89 | 0.84 | 0.89 | 0.82 | 0.83 | 0.77 | 0.79 | 0.61 | 0.62 | 0.61 | 0.68 |
| Prune Head 3 | 0.06 | 0.91 | 0.92 | 0.9 | 0.08 | 0.71 | 0.61 | 0.71 | 0.07 | 0.56 | 0.61 | 0.49 | 0.07 | 0.6 | 0.59 | 0.6 | 0.1 | 0.85 | 0.89 | 0.90 | 0.12 | 0.79 | 0.83 | 0.79 | 0.09 | 0.62 | 0.66 | 0.68 |
| Prune Head 4 | 0.89 | 0.92 | 0.93 | 0.9 | 0.71 | 0.66 | 0.66 | 0.68 | 0.59 | 0.54 | 0.54 | 0.51 | 0.59 | 0.6 | 0.55 | 0.59 | 0.85 | 0.82 | 0.87 | 0.89 | 0.82 | 0.78 | 0.77 | 0.78 | 0.62 | 0.62 | 0.62 | 0.68 |
| Prune Head 5 | 0.56 | 0.91 | 0.9 | 0.91 | 0.5 | 0.61 | 0.61 | 0.66 | 0.46 | 0.59 | 0.54 | 0.56 | 0.37 | 0.58 | 0.53 | 0.59 | 0.47 | 0.89 | 0.81 | 0.84 | 0.54 | 0.83 | 0.76 | 0.77 | 0.41 | 0.64 | 0.68 | 0.66 |
| Prune Head 6 | 0.81 | 0.92 | 0.9 | 0.9 | 0.61 | 0.66 | 0.63 | 0.63 | 0.56 | 0.56 | 0.59 | 0.51 | 0.51 | 0.59 | 0.59 | 0.57 | 0.71 | 0.82 | 0.82 | 0.89 | 0.73 | 0.82 | 0.74 | 0.76 | 0.54 | 0.66 | 0.66 | 0.64 |
| Prune Head 7 | 0.91 | 0.92 | 0.9 | 0.9 | 0.63 | 0.61 | 0.63 | 0.66 | 0.56 | 0.56 | 0.54 | 0.54 | 0.58 | 0.6 | 0.56 | 0.59 | 0.85 | 0.87 | 0.89 | 0.87 | 0.78 | 0.79 | 0.8 | 0.78 | 0.64 | 0.68 | 0.55 | 0.62 |
| Prune Head 8 | 0.86 | 0.93 | 0.9 | 0.93 | 0.63 | 0.74 | 0.68 | 0.68 | 0.59 | 0.49 | 0.56 | 0.54 | 0.53 | 0.56 | 0.59 | 0.58 | 0.82 | 0.85 | 0.87 | 0.85 | 0.79 | 0.8 | 0.82 | 0.80 | 0.62 | 0.61 | 0.62 | 0.64 |
| Prune Head 9 | 0.93 | 0.91 | 0.91 | 0.91 | 0.66 | 0.66 | 0.68 | 0.66 | 0.56 | 0.49 | 0.54 | 0.49 | 0.57 | 0.54 | 0.52 | 0.57 | 0.9 | 0.84 | 0.85 | 0.89 | 0.78 | 0.82 | 0.8 | 0.79 | 0.64 | 0.66 | 0.66 | 0.66 |
| Prune Head 10 | 0.92 | 0.92 | 0.92 | 0.89 | 0.68 | 0.66 | 0.68 | 0.66 | 0.54 | 0.51 | 0.56 | 0.51 | 0.55 | 0.57 | 0.56 | 0.57 | 0.85 | 0.85 | 0.84 | 0.87 | 0.8 | 0.8 | 0.77 | 0.78 | 0.66 | 0.62 | 0.62 | 0.70 |
| Prune Head 11 | 0.9 | 0.92 | 0.91 | 0.9 | 0.68 | 0.66 | 0.71 | 0.66 | 0.51 | 0.56 | 0.56 | 0.49 | 0.6 | 0.61 | 0.55 | 0.6 | 0.87 | 0.84 | 0.87 | 0.87 | 0.78 | 0.8 | 0.77 | 0.80 | 0.66 | 0.62 | 0.61 | 0.66 |
| No Prune | 0.91 | 0.91 | 0.91 | 0.91 | 0.63 | 0.63 | 0.63 | 0.63 | 0.49 | 0.49 | 0.49 | 0.49 | 0.58 | 0.58 | 0.58 | 0.57 | 0.89 | 0.89 | 0.89 | 0.87 | 0.79 | 0.79 | 0.79 | 0.79 | 0.66 | 0.66 | 0.66 | 0.68 |
| Algebra | Counting & Probability | Geometry | Intermediate Algebra | Number Theory | Prealgebra | Precalculus | ||||||||||||||||||||||
| Layer | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 |
| Prune Head 0 | 0.96 | 0.96 | 0.97 | 0.97 | 0.84 | 0.82 | 0.84 | 0.84 | 0.68 | 0.68 | 0.59 | 0.66 | 0.68 | 0.67 | 0.64 | 0.62 | 0.9 | 0.95 | 0.95 | 0.97 | 0.9 | 0.88 | 0.9 | 0.88 | 0.71 | 0.77 | 0.7 | 0.80 |
| Prune Head 1 | 0.97 | 0.96 | 0.97 | 0.97 | 0.84 | 0.79 | 0.84 | 0.82 | 0.68 | 0.66 | 0.66 | 0.66 | 0.63 | 0.65 | 0.63 | 0.62 | 0.95 | 0.95 | 0.97 | 0.95 | 0.89 | 0.89 | 0.89 | 0.88 | 0.73 | 0.75 | 0.75 | 0.77 |
| Prune Head 2 | 0.97 | 0.96 | 0.98 | 0.97 | 0.84 | 0.84 | 0.87 | 0.84 | 0.66 | 0.66 | 0.66 | 0.66 | 0.65 | 0.64 | 0.64 | 0.65 | 0.92 | 0.95 | 0.92 | 0.95 | 0.9 | 0.89 | 0.9 | 0.87 | 0.75 | 0.73 | 0.71 | 0.77 |
| Prune Head 3 | 0.36 | 0.97 | 0.97 | 0.97 | 0.5 | 0.82 | 0.82 | 0.82 | 0.29 | 0.61 | 0.61 | 0.66 | 0.25 | 0.65 | 0.68 | 0.62 | 0.29 | 0.97 | 1.00 | 0.95 | 0.34 | 0.89 | 0.88 | 0.88 | 0.2 | 0.77 | 0.73 | 0.75 |
| Prune Head 4 | 0.97 | 0.97 | 0.95 | 0.97 | 0.84 | 0.82 | 0.82 | 0.82 | 0.68 | 0.66 | 0.68 | 0.66 | 0.61 | 0.66 | 0.63 | 0.62 | 0.95 | 0.94 | 0.92 | 0.95 | 0.89 | 0.90 | 0.89 | 0.90 | 0.75 | 0.73 | 0.75 | 0.75 |
| Prune Head 5 | 0.96 | 0.95 | 0.96 | 0.97 | 0.82 | 0.82 | 0.84 | 0.82 | 0.73 | 0.63 | 0.63 | 0.68 | 0.63 | 0.67 | 0.65 | 0.62 | 0.95 | 0.94 | 0.97 | 0.92 | 0.89 | 0.90 | 0.88 | 0.88 | 0.73 | 0.75 | 0.71 | 0.75 |
| Prune Head 6 | 0.97 | 0.96 | 0.96 | 0.97 | 0.84 | 0.82 | 0.82 | 0.82 | 0.63 | 0.68 | 0.63 | 0.66 | 0.66 | 0.61 | 0.67 | 0.63 | 0.92 | 0.94 | 0.95 | 0.95 | 0.88 | 0.89 | 0.9 | 0.89 | 0.68 | 0.73 | 0.77 | 0.79 |
| Prune Head 7 | 0.96 | 0.97 | 0.97 | 0.97 | 0.79 | 0.82 | 0.82 | 0.82 | 0.61 | 0.66 | 0.63 | 0.73 | 0.64 | 0.64 | 0.66 | 0.62 | 0.95 | 0.95 | 0.95 | 0.95 | 0.89 | 0.89 | 0.89 | 0.89 | 0.73 | 0.79 | 0.75 | 0.71 |
| Prune Head 8 | 0.97 | 0.97 | 0.97 | 0.97 | 0.82 | 0.84 | 0.84 | 0.84 | 0.68 | 0.66 | 0.68 | 0.68 | 0.65 | 0.63 | 0.65 | 0.63 | 0.97 | 0.95 | 0.94 | 0.97 | 0.89 | 0.88 | 0.88 | 0.88 | 0.75 | 0.75 | 0.7 | 0.71 |
| Prune Head 9 | 0.96 | 0.97 | 0.96 | 0.97 | 0.84 | 0.84 | 0.82 | 0.84 | 0.61 | 0.66 | 0.59 | 0.76 | 0.63 | 0.65 | 0.64 | 0.67 | 0.94 | 0.95 | 0.97 | 0.95 | 0.89 | 0.88 | 0.9 | 0.89 | 0.77 | 0.75 | 0.77 | 0.73 |
| Prune Head 10 | 0.95 | 0.97 | 0.96 | 0.96 | 0.84 | 0.82 | 0.84 | 0.84 | 0.66 | 0.63 | 0.63 | 0.71 | 0.64 | 0.63 | 0.65 | 0.64 | 0.94 | 0.95 | 0.94 | 0.95 | 0.89 | 0.88 | 0.89 | 0.88 | 0.73 | 0.75 | 0.71 | 0.75 |
| Prune Head 11 | 0.97 | 0.97 | 0.97 | 0.97 | 0.82 | 0.82 | 0.82 | 0.84 | 0.63 | 0.71 | 0.63 | 0.73 | 0.64 | 0.65 | 0.69 | 0.63 | 0.97 | 0.94 | 0.95 | 0.95 | 0.9 | 0.88 | 0.9 | 0.88 | 0.79 | 0.71 | 0.77 | 0.75 |
| Prune Head 12 | 0.97 | 0.97 | 0.98 | 0.97 | 0.82 | 0.82 | 0.87 | 0.84 | 0.66 | 0.66 | 0.68 | 0.71 | 0.66 | 0.63 | 0.67 | 0.64 | 0.95 | 0.95 | 0.92 | 0.97 | 0.88 | 0.89 | 0.87 | 0.88 | 0.77 | 0.77 | 0.75 | 0.73 |
| Prune Head 13 | 0.97 | 0.97 | 0.98 | 0.97 | 0.82 | 0.82 | 0.82 | 0.84 | 0.61 | 0.68 | 0.68 | 0.68 | 0.6 | 0.64 | 0.63 | 0.64 | 0.95 | 0.95 | 0.94 | 0.97 | 0.88 | 0.88 | 0.91 | 0.88 | 0.73 | 0.75 | 0.79 | 0.75 |
| Prune Head 14 | 0.96 | 0.97 | 0.96 | 0.97 | 0.84 | 0.82 | 0.84 | 0.82 | 0.63 | 0.61 | 0.66 | 0.66 | 0.67 | 0.66 | 0.63 | 0.61 | 0.94 | 0.95 | 0.95 | 0.95 | 0.89 | 0.88 | 0.89 | 0.85 | 0.77 | 0.73 | 0.71 | 0.75 |
| Prune Head 15 | 0.97 | 0.97 | 0.94 | 0.97 | 0.79 | 0.82 | 0.84 | 0.82 | 0.73 | 0.68 | 0.63 | 0.61 | 0.66 | 0.66 | 0.64 | 0.63 | 0.97 | 0.94 | 0.94 | 0.95 | 0.9 | 0.88 | 0.91 | 0.88 | 0.77 | 0.71 | 0.75 | 0.75 |
| Prune Head 16 | 0.96 | 0.96 | 0.97 | 0.97 | 0.82 | 0.84 | 0.82 | 0.84 | 0.68 | 0.63 | 0.68 | 0.66 | 0.67 | 0.62 | 0.64 | 0.63 | 0.9 | 0.95 | 0.97 | 0.95 | 0.88 | 0.89 | 0.89 | 0.87 | 0.73 | 0.75 | 0.75 | 0.75 |
| Prune Head 17 | 0.96 | 0.96 | 0.98 | 0.97 | 0.82 | 0.82 | 0.82 | 0.82 | 0.66 | 0.63 | 0.68 | 0.61 | 0.65 | 0.66 | 0.64 | 0.63 | 0.92 | 0.94 | 0.92 | 0.95 | 0.87 | 0.88 | 0.89 | 0.88 | 0.75 | 0.75 | 0.75 | 0.75 |
| Prune Head 18 | 0.96 | 0.97 | 0.97 | 0.97 | 0.82 | 0.82 | 0.87 | 0.82 | 0.61 | 0.63 | 0.66 | 0.66 | 0.67 | 0.6 | 0.61 | 0.63 | 0.94 | 0.94 | 0.95 | 0.95 | 0.89 | 0.89 | 0.88 | 0.88 | 0.77 | 0.77 | 0.77 | 0.75 |
| Prune Head 19 | 0.96 | 0.96 | 0.97 | 0.97 | 0.82 | 0.82 | 0.79 | 0.82 | 0.63 | 0.68 | 0.66 | 0.63 | 0.67 | 0.64 | 0.65 | 0.64 | 0.92 | 0.95 | 0.95 | 0.95 | 0.87 | 0.89 | 0.89 | 0.90 | 0.73 | 0.71 | 0.71 | 0.75 |
| Prune Head 20 | 0.96 | 0.97 | 0.96 | 0.97 | 0.84 | 0.82 | 0.82 | 0.82 | 0.63 | 0.61 | 0.66 | 0.68 | 0.65 | 0.62 | 0.65 | 0.62 | 0.92 | 0.95 | 0.94 | 0.95 | 0.88 | 0.90 | 0.89 | 0.88 | 0.75 | 0.71 | 0.73 | 0.75 |
| Prune Head 21 | 0.97 | 0.96 | 0.97 | 0.97 | 0.79 | 0.84 | 0.82 | 0.84 | 0.63 | 0.59 | 0.66 | 0.66 | 0.64 | 0.62 | 0.66 | 0.61 | 0.9 | 0.95 | 0.95 | 0.95 | 0.91 | 0.88 | 0.89 | 0.88 | 0.73 | 0.75 | 0.71 | 0.71 |
| Prune Head 22 | 0.97 | 0.97 | 0.96 | 0.97 | 0.84 | 0.84 | 0.82 | 0.84 | 0.71 | 0.66 | 0.68 | 0.66 | 0.62 | 0.63 | 0.62 | 0.61 | 0.95 | 0.95 | 0.97 | 0.94 | 0.9 | 0.89 | 0.88 | 0.88 | 0.73 | 0.73 | 0.75 | 0.75 |
| Prune Head 23 | 0.96 | 0.97 | 0.98 | 0.97 | 0.84 | 0.84 | 0.82 | 0.82 | 0.63 | 0.66 | 0.61 | 0.63 | 0.65 | 0.64 | 0.65 | 0.64 | 0.95 | 0.94 | 0.94 | 0.95 | 0.9 | 0.89 | 0.9 | 0.88 | 0.73 | 0.73 | 0.77 | 0.75 |
| Prune Head 24 | 0.97 | 0.96 | 0.96 | 0.97 | 0.82 | 0.82 | 0.79 | 0.84 | 0.73 | 0.63 | 0.66 | 0.63 | 0.65 | 0.62 | 0.66 | 0.62 | 0.94 | 0.95 | 0.9 | 0.95 | 0.89 | 0.88 | 0.9 | 0.89 | 0.73 | 0.77 | 0.77 | 0.77 |
| Prune Head 25 | 0.97 | 0.97 | 0.96 | 0.97 | 0.84 | 0.82 | 0.79 | 0.82 | 0.63 | 0.71 | 0.68 | 0.66 | 0.64 | 0.68 | 0.61 | 0.61 | 0.95 | 0.94 | 0.94 | 0.95 | 0.88 | 0.89 | 0.91 | 0.90 | 0.75 | 0.73 | 0.73 | 0.77 |
| Prune Head 26 | 0.97 | 0.98 | 0.98 | 0.97 | 0.79 | 0.82 | 0.79 | 0.84 | 0.71 | 0.68 | 0.68 | 0.68 | 0.62 | 0.67 | 0.62 | 0.61 | 0.95 | 0.95 | 0.95 | 0.94 | 0.88 | 0.88 | 0.88 | 0.89 | 0.73 | 0.79 | 0.75 | 0.77 |
| Prune Head 27 | 0.97 | 0.97 | 0.97 | 0.97 | 0.84 | 0.79 | 0.82 | 0.82 | 0.66 | 0.66 | 0.66 | 0.61 | 0.62 | 0.65 | 0.64 | 0.63 | 0.95 | 0.95 | 0.95 | 0.94 | 0.88 | 0.90 | 0.91 | 0.88 | 0.73 | 0.73 | 0.68 | 0.75 |
| No Prune | 0.97 | 0.97 | 0.97 | 0.97 | 0.82 | 0.82 | 0.82 | 0.82 | 0.66 | 0.66 | 0.66 | 0.66 | 0.62 | 0.62 | 0.62 | 0.62 | 0.95 | 0.95 | 0.95 | 0.95 | 0.88 | 0.88 | 0.88 | 0.88 | 0.75 | 0.75 | 0.75 | 0.75 |
| Algebra | Counting & Probability | Geometry | Intermediate Algebra | Number Theory | Prealgebra | Precalculus | ||||||||||||||||||||||
| Layer | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 | 0 | 5 | 15 | 27 |
| Prune Head 0 | 0.55 | 0.94 | 0.95 | 0.95 | 0.5 | 0.74 | 0.76 | 0.71 | 0.44 | 0.56 | 0.56 | 0.59 | 0.38 | 0.55 | 0.54 | 0.57 | 0.34 | 0.85 | 0.85 | 0.87 | 0.49 | 0.83 | 0.84 | 0.87 | 0.36 | 0.59 | 0.61 | 0.59 |
| Prune Head 1 | 0.85 | 0.94 | 0.93 | 0.94 | 0.61 | 0.71 | 0.74 | 0.71 | 0.49 | 0.63 | 0.56 | 0.61 | 0.51 | 0.57 | 0.53 | 0.59 | 0.82 | 0.87 | 0.87 | 0.87 | 0.73 | 0.84 | 0.83 | 0.87 | 0.70 | 0.62 | 0.59 | 0.61 |
| Prune Head 2 | 0.94 | 0.94 | 0.95 | 0.94 | 0.68 | 0.71 | 0.71 | 0.71 | 0.63 | 0.61 | 0.71 | 0.66 | 0.58 | 0.56 | 0.54 | 0.58 | 0.82 | 0.84 | 0.81 | 0.85 | 0.87 | 0.84 | 0.83 | 0.87 | 0.61 | 0.61 | 0.61 | 0.64 |
| Prune Head 3 | 0.02 | 0.94 | 0.92 | 0.94 | 0 | 0.76 | 0.76 | 0.74 | 0.05 | 0.61 | 0.68 | 0.61 | 0.04 | 0.58 | 0.60 | 0.6 | 0 | 0.85 | 0.84 | 0.85 | 0.01 | 0.84 | 0.84 | 0.87 | 0.02 | 0.61 | 0.62 | 0.61 |
| Prune Head 4 | 0.94 | 0.94 | 0.94 | 0.94 | 0.71 | 0.68 | 0.74 | 0.71 | 0.61 | 0.59 | 0.56 | 0.61 | 0.59 | 0.58 | 0.54 | 0.59 | 0.85 | 0.82 | 0.79 | 0.85 | 0.87 | 0.84 | 0.83 | 0.85 | 0.62 | 0.57 | 0.64 | 0.61 |
| Prune Head 5 | 0.95 | 0.95 | 0.93 | 0.94 | 0.71 | 0.74 | 0.74 | 0.71 | 0.59 | 0.61 | 0.61 | 0.56 | 0.58 | 0.54 | 0.57 | 0.6 | 0.85 | 0.84 | 0.84 | 0.82 | 0.87 | 0.82 | 0.84 | 0.85 | 0.62 | 0.62 | 0.62 | 0.61 |
| Prune Head 6 | 0.63 | 0.95 | 0.91 | 0.95 | 0.53 | 0.76 | 0.66 | 0.71 | 0.54 | 0.59 | 0.56 | 0.61 | 0.44 | 0.57 | 0.53 | 0.6 | 0.45 | 0.89 | 0.84 | 0.87 | 0.65 | 0.85 | 0.83 | 0.85 | 0.39 | 0.62 | 0.64 | 0.62 |
| Prune Head 7 | 0.96 | 0.93 | 0.91 | 0.94 | 0.76 | 0.74 | 0.76 | 0.74 | 0.68 | 0.61 | 0.54 | 0.59 | 0.55 | 0.57 | 0.59 | 0.58 | 0.84 | 0.87 | 0.82 | 0.87 | 0.84 | 0.85 | 0.85 | 0.84 | 0.64 | 0.64 | 0.61 | 0.64 |
| Prune Head 8 | 0.94 | 0.95 | 0.94 | 0.94 | 0.74 | 0.74 | 0.74 | 0.74 | 0.63 | 0.63 | 0.59 | 0.61 | 0.58 | 0.58 | 0.58 | 0.58 | 0.82 | 0.87 | 0.84 | 0.84 | 0.85 | 0.85 | 0.85 | 0.85 | 0.55 | 0.57 | 0.59 | 0.59 |
| Prune Head 9 | 0.95 | 0.95 | 0.94 | 0.97 | 0.68 | 0.74 | 0.68 | 0.71 | 0.56 | 0.66 | 0.56 | 0.59 | 0.58 | 0.55 | 0.57 | 0.57 | 0.85 | 0.85 | 0.92 | 0.87 | 0.87 | 0.87 | 0.87 | 0.85 | 0.61 | 0.57 | 0.62 | 0.64 |
| Prune Head 10 | 0.91 | 0.94 | 0.94 | 0.94 | 0.61 | 0.74 | 0.68 | 0.76 | 0.54 | 0.61 | 0.63 | 0.61 | 0.60 | 0.6 | 0.55 | 0.56 | 0.82 | 0.85 | 0.77 | 0.85 | 0.82 | 0.85 | 0.8 | 0.8 | 0.61 | 0.61 | 0.64 | 0.55 |
| Prune Head 11 | 0.87 | 0.93 | 0.95 | 0.94 | 0.71 | 0.74 | 0.76 | 0.74 | 0.59 | 0.61 | 0.56 | 0.61 | 0.57 | 0.56 | 0.55 | 0.57 | 0.77 | 0.87 | 0.89 | 0.82 | 0.83 | 0.84 | 0.84 | 0.84 | 0.52 | 0.62 | 0.68 | 0.61 |
| Prune Head 12 | 0.94 | 0.94 | 0.94 | 0.97 | 0.74 | 0.71 | 0.74 | 0.74 | 0.63 | 0.59 | 0.59 | 0.61 | 0.56 | 0.59 | 0.57 | 0.57 | 0.84 | 0.85 | 0.89 | 0.87 | 0.87 | 0.84 | 0.82 | 0.82 | 0.68 | 0.61 | 0.57 | 0.57 |
| Prune Head 13 | 0.95 | 0.95 | 0.93 | 0.94 | 0.71 | 0.71 | 0.71 | 0.79 | 0.61 | 0.63 | 0.61 | 0.59 | 0.57 | 0.6 | 0.60 | 0.54 | 0.85 | 0.89 | 0.87 | 0.85 | 0.87 | 0.88 | 0.83 | 0.83 | 0.62 | 0.59 | 0.55 | 0.55 |
| Prune Head 14 | 0.94 | 0.93 | 0.94 | 0.94 | 0.71 | 0.71 | 0.68 | 0.76 | 0.63 | 0.56 | 0.61 | 0.59 | 0.54 | 0.55 | 0.55 | 0.58 | 0.77 | 0.84 | 0.85 | 0.84 | 0.82 | 0.85 | 0.84 | 0.87 | 0.70 | 0.61 | 0.64 | 0.61 |
| Prune Head 15 | 0.81 | 0.94 | 0.94 | 0.95 | 0.68 | 0.74 | 0.76 | 0.71 | 0.54 | 0.56 | 0.61 | 0.61 | 0.52 | 0.56 | 0.58 | 0.59 | 0.77 | 0.84 | 0.85 | 0.82 | 0.82 | 0.88 | 0.85 | 0.88 | 0.59 | 0.57 | 0.64 | 0.59 |
| Prune Head 16 | 0.96 | 0.94 | 0.94 | 0.94 | 0.68 | 0.68 | 0.76 | 0.74 | 0.66 | 0.61 | 0.63 | 0.59 | 0.58 | 0.61 | 0.54 | 0.59 | 0.84 | 0.82 | 0.87 | 0.84 | 0.88 | 0.87 | 0.82 | 0.88 | 0.59 | 0.62 | 0.62 | 0.62 |
| Prune Head 17 | 0.93 | 0.95 | 0.94 | 0.94 | 0.71 | 0.71 | 0.74 | 0.76 | 0.59 | 0.59 | 0.63 | 0.59 | 0.54 | 0.55 | 0.60 | 0.61 | 0.81 | 0.84 | 0.87 | 0.85 | 0.89 | 0.87 | 0.82 | 0.82 | 0.62 | 0.59 | 0.61 | 0.61 |
| Prune Head 18 | 0.94 | 0.94 | 0.94 | 0.94 | 0.74 | 0.71 | 0.74 | 0.71 | 0.66 | 0.61 | 0.56 | 0.61 | 0.56 | 0.56 | 0.58 | 0.56 | 0.85 | 0.84 | 0.85 | 0.87 | 0.87 | 0.85 | 0.83 | 0.84 | 0.64 | 0.59 | 0.66 | 0.61 |
| Prune Head 19 | 0.94 | 0.96 | 0.92 | 0.94 | 0.71 | 0.76 | 0.76 | 0.71 | 0.63 | 0.56 | 0.66 | 0.59 | 0.56 | 0.55 | 0.53 | 0.58 | 0.82 | 0.87 | 0.84 | 0.85 | 0.87 | 0.79 | 0.87 | 0.85 | 0.66 | 0.57 | 0.70 | 0.61 |
| Prune Head 20 | 0.95 | 0.94 | 0.94 | 0.94 | 0.74 | 0.74 | 0.74 | 0.71 | 0.66 | 0.59 | 0.59 | 0.61 | 0.60 | 0.57 | 0.57 | 0.58 | 0.81 | 0.87 | 0.82 | 0.84 | 0.84 | 0.82 | 0.84 | 0.88 | 0.61 | 0.62 | 0.66 | 0.61 |
| Prune Head 21 | 0.95 | 0.96 | 0.96 | 0.96 | 0.76 | 0.74 | 0.74 | 0.71 | 0.61 | 0.61 | 0.61 | 0.61 | 0.56 | 0.58 | 0.58 | 0.6 | 0.85 | 0.85 | 0.85 | 0.87 | 0.85 | 0.85 | 0.84 | 0.87 | 0.61 | 0.62 | 0.62 | 0.61 |
| Prune Head 22 | 0.44 | 0.96 | 0.93 | 0.94 | 0.24 | 0.76 | 0.74 | 0.71 | 0.22 | 0.59 | 0.66 | 0.59 | 0.12 | 0.58 | 0.55 | 0.61 | 0.26 | 0.85 | 0.87 | 0.84 | 0.52 | 0.89 | 0.83 | 0.84 | 0.07 | 0.62 | 0.64 | 0.61 |
| Prune Head 23 | 0.11 | 0.94 | 0.92 | 0.94 | 0.21 | 0.74 | 0.63 | 0.71 | 0.17 | 0.66 | 0.68 | 0.59 | 0.04 | 0.62 | 0.49 | 0.58 | 0.06 | 0.84 | 0.79 | 0.87 | 0.27 | 0.85 | 0.82 | 0.87 | 0.04 | 0.61 | 0.59 | 0.59 |
| Prune Head 24 | 0.83 | 0.94 | 0.93 | 0.96 | 0.61 | 0.74 | 0.71 | 0.71 | 0.56 | 0.66 | 0.63 | 0.66 | 0.51 | 0.6 | 0.56 | 0.6 | 0.77 | 0.85 | 0.84 | 0.85 | 0.68 | 0.87 | 0.84 | 0.84 | 0.57 | 0.62 | 0.55 | 0.62 |
| Prune Head 25 | 0.02 | 0.95 | 0.95 | 0.95 | 0 | 0.76 | 0.74 | 0.66 | 0 | 0.63 | 0.66 | 0.63 | 0 | 0.55 | 0.60 | 0.59 | 0 | 0.89 | 0.81 | 0.85 | 0.02 | 0.84 | 0.83 | 0.85 | 0.02 | 0.70 | 0.59 | 0.66 |
| Prune Head 26 | 0.5 | 0.95 | 0.93 | 0.95 | 0.47 | 0.71 | 0.71 | 0.71 | 0.37 | 0.54 | 0.54 | 0.56 | 0.25 | 0.58 | 0.60 | 0.56 | 0.56 | 0.85 | 0.89 | 0.89 | 0.63 | 0.88 | 0.85 | 0.85 | 0.21 | 0.55 | 0.64 | 0.66 |
| Prune Head 27 | 0.83 | 0.96 | 0.94 | 0.95 | 0.68 | 0.71 | 0.71 | 0.63 | 0.46 | 0.68 | 0.59 | 0.56 | 0.49 | 0.6 | 0.52 | 0.59 | 0.69 | 0.85 | 0.82 | 0.82 | 0.63 | 0.84 | 0.84 | 0.85 | 0.61 | 0.59 | 0.59 | 0.64 |
| No Prune | 0.94 | 0.94 | 0.94 | 0.94 | 0.74 | 0.74 | 0.74 | 0.74 | 0.59 | 0.59 | 0.59 | 0.59 | 0.59 | 0.59 | 0.59 | 0.59 | 0.84 | 0.84 | 0.84 | 0.84 | 0.87 | 0.87 | 0.87 | 0.87 | 0.59 | 0.59 | 0.59 | 0.59 |